# Semantic Exploration and Dense Mapping of Complex Environments using Ground Robots Equipped with LiDAR and Panoramic Camera

### v1

@@ Abstract
This paper presents a system for autonomous semantic exploration and dense semantic target mapping of a complex unknown environment using a ground robot equipped with a LiDAR-panoramic camera suite. Existing approaches often struggle to strike a balance between collecting enough high-quality observations from multiple view angles and avoiding unnecessary repetitive traversal. To fill this gap, we propose a complete system that combines mapping and planning. We first redefine the task as completing both geometric coverage and semantic viewpoint observation. Subsequently, we manage semantic and geometric viewpoints separately and propose a novel Priority-driven Decoupled Local Sampler to generate local viewpoint sets. This allows for explicit multi-view semantic inspection and voxel coverage without unnecessary repetition. Building on this, we develop a hierarchical planner that ensures efficient global coverage. In addition, we propose a Safe Aggressive Exploration State Machine, which allows aggressive exploration behavior while ensuring the robot’s safety. Our system includes a plug-and-play semantic target mapping module that integrates seamlessly with state-of-the-art SLAM algorithms for pointcloud-level dense semantic target mapping. We validate our approach through extensive experiments in both realistic simulations and complex real-world environments. Simulation results demonstrate that our planner achieves faster exploration and shorter travel distances while guaranteeing a specified number of multi-view inspections. Real-world experiments further confirm the effectiveness of the system in achieving accurate dense semantic object mapping of unstructured environments.
@@ I INTRODUCTION
With the growing success of autonomous exploration algorithms, their applications are drawing increased attention from the research community. Semantic exploration is particularly important in scenarios such as factories and construction sites. Unlike unmanned aerial vehicles (UAVs) commonly used in prior work, ground robots offer the advantage of accommodating more powerful sensors and computing systems, including high-resolution LiDAR and panoramic camera arrays. This paper focuses on semantic exploration and mapping for ground robots.
@@
Despite advancements, autonomous semantic exploration faces challenges. First, it requires an effective integration of mapping and planning, as the planner depends on semantic information to make informed decisions. To achieve high-quality mapping, the planner needs to collect comprehensive measurements from multiple distinct viewpoints while computing a path for geometric coverage that is as short as possible. Existing methods either assign semantic weights to potential exploration goals that prioritize positions near target objects [1][2] or sample viewpoints that cover unknown regions [3]. However, the former often causes repetitive movements around objects, while the latter neglects the importance of complete and diverse observation from multiple view angles for target objects. There is still a lack of a planner that strikes the balance for both. Another difficulty is that robots operating in complicated environments need to balance exploration speed and collision risks. Existing planners often limit exploration to known free spaces for safety, but this over-conservative strategy results in short-sighted decisions and inefficient paths [4]. However, aggressive exploration in unknown areas raises collision risks, especially at higher speeds, which limits planner performance.
@@
To address the above challenges, this paper proposes a system that performs semantic exploration planning and dense semantic target mapping in an unknown environment. We develop a modular, object-centric semantic target dense mapping component that works seamlessly with various state-of-the-art SLAM systems. The planning component defines the semantic exploration task as inspecting all unknown voxels and visiting evenly distributed semantic viewpoints to ensure high-quality mapping. To support the integration between mapping and planning, we introduce a hierarchical planner that manages semantic viewpoints and geometric exploration viewpoints in a unified framework. We propose a novel Priority-driven Decoupled Local Sampler to handle the two types of viewpoints separately and generate a minimum set of viewpoints that supports the completeness of both semantic and geometric coverage tasks. The sampler is then integrated with a global planner to generate the global optimal path by solving an asymmetric traveling salesman problem (ATSP). After that, to support aggressive exploration, a novel Safe Aggressive Exploration State Machine is proposed to ensure aggressive and resilient exploration in crowded regions.
@@
  • A plug-in semantic target mapping module based on LiDAR-Panoramic camera fusion, tailored to the semantic exploration planner. This module can be integrated with any state-of-the-art SLAM system.
@@
  • An aggressive exploration state machine that supports the robot in choosing aggressive goals, maintaining safety in unknown space, and recovering from the near collision state.
  • A complete ground robot system with a specialized sensor suite, modular mapping pipeline, and a novel planning framework. Simulated and real-world tests in industrial and construction environments validate the effectiveness of the system.
@@ II RELATED WORK
Autonomous robot exploration has made great progress over the decades [5][6][7][8]. One stream of methods is based on frontiers, which are the boundaries between free space and unknown space. Yamauchi [5] greedily chooses the closest frontier as the next goal, while later methods [6][9] apply more sophisticated frontier clustering, viewpoint generation, and path planning techniques. Another stream of methods is based on Next-Best-View (NBV) and graph sampling, such as RRT or PRM. In [10], a graph is built by sampling in safe free space and calculating the information gain for each node. The next best node with the highest average gain is chosen as the next goal. Later, Dang et al. [8] propose rapidly-exploring random graphs for more rapid NBV exploration planning, and other methods [11][12][13][14][15] continue to propose more carefully designed utility functions for information gain and graph structures. In recent years, more algorithms have been formulated the traveling salesman problem (TSP)[6][7][16][3] to improve the efficiency of traversing all viewpoints or graph nodes.
@@
At the same time, the SLAM community has an increasing interest in metric-semantic SLAM. Some of them use sparse representations such as cubes [17], ellipsoids [18], or combinations of geometric shapes [19], and combine the parameters into a factor graph to formulate an optimization problem. Others use dense representations such as meshes [20] or sufels [21].
@@
With the development of both areas, exploration with semantic objects is gaining more attention. Tung et al. [2] define an object gain function to handle re-observation from multiple resolutions, and Sotiris et al. [1] propose multiple utility functions to include both background and object information gain to improve the quality of camera-based mesh reconstruction. Yiming et al. [3] further propose a hierarchical planning framework to solve exploration with target search. At the same time, the DARPA subterranean challenge facilitated semantic exploration research. Xianmei et al. [22] propose a whole pipeline to guarantee high precision and recall of semantic detection during exploration and emphasize the importance of observation from multiple view angles. Graeme et al. [23] utilized the fusion of multiple sensors during exploration and mapping.
@@
None of these methods, however, explicitly defines multi-view observation viewpoints for object mapping in exploration, thus suffering from randomness due to unnecessary re-observation or inadequate observation. In addition, most algorithms are still based on the greedy NBV strategy with semantic weights, lacking careful consideration of global optimality. Our method fills this gap by explicitly defining the semantic viewpoint with a certain density around target objects while enabling efficient traversing considering both geometric coverage and semantic observation. Furthermore, we developed a Safe Aggressive State Machine to support aggressive exploration in complicated real-world environments.
@@ III PROBLEM FORMULATION
The goal of this paper is to enable a single ground robot, equipped with a LiDAR and panoramic camera fusion system, to autonomously plan its trajectory within an unknown environment and accurately map both the environment itself and specified target objects of interest. The robot should generate a dense point cloud for both the environment and the identified objects, while optimizing trajectory efficiency to ensure timely completion of the mapping tasks.
@@ III-A Robot and Sensor Model
In this work, we employ the Boston Dynamics Spot as a quadruped mobile platform equipped with a specialized sensor system. The sensor suite includes an Ouster OS-1-128 LiDAR and a panoramic camera system. The LiDAR features a 45°vertical field of view (FOV) and a 360°horizontal FOV, with 128 vertical channels, operating at a frequency of 10 Hz. The panoramic camera system is the Boston Dynamics CAM Payload, which integrates data from five Sony IMX290 cameras to provide a 360°$\times$ 165°field of view.
@@ III-B Environment Model
The unknown environment is modeled as a bounded 3D volume $V\subset\mathbb{R}^{3}$, discretized into individual voxels $v$. Each voxel $v$ is associated with an occupancy probability $P_{o}(v)$. Before receiving any sensor measurements, each voxel is initialized as unknown with $P_{o}(v)=0.5$. Through the accumulation of sensor measurements and subsequent ray-casting, voxels are eventually classified as either occupied $V_{occ}$ or free $V_{free}$. However, due to the complexity of environment geometry, robot configuration constraints, and sensor limitations, some residual voxels $V_{res}$ may remain unmapped. Hence, the relationship $V_{free}\cup V_{occ}\cup V_{res}=V$ holds at the end of the exploration.
@@ III-C Problem Definition
Given the robot, sensor, and environment models, the overall problem can be defined as:
Problem 1 (Semantic Target Mapping): Assume an unknown number $N\subset\mathbf{N}$ of target objects, each associated with a semantic label $\lambda$ from a semantic label set $\mathbb{L}$. The robot must detect and classify all target objects, extracting the dense point cloud for both the target objects and the surrounding environment.
As discussed in Section I, the semantic exploration task must consider both target object observation and geometric coverage. We divide the space centering around a target object $O_{i}$ with radius $r$ into $M$ evenly distributed sectors. Each sector $S^{i}_{m}$ is defined as an observation space. Then, we define the semantic exploration task as follows:
Problem 2 (Semantic Exploration): Given the sensor model and configuration $\psi_{robot}$, plan a path $\sigma_{robot}$ starting from an initial configuration $\xi_{robot}$ such that: (1) it covers at least one valid detection in every observation space $S^{i}_{m}$ for any target object $O_{i}$, whenever possible; and (2) it observes all voxels so that $V_{free}\cup V_{occ}=V\setminus V_{res}$.
@@ IV METHODOLOGY
The proposed system comprises a plug-in semantic mapping module and a hierarchical planner, as shown in Fig. 2. The mapping module, discussed in Section IV-B, takes as input the panoramic image $\mathbf{I}$, LiDAR point cloud $\mathbf{P}$, and odometry $\mathbf{O}$, then outputs the dense point cloud map of both background environment $\mathbf{M_{b}}$ and target objects $\mathbf{M_{o}}.$ Concurrently, the hierarchical planner generates the exploration path that ensures both semantic observation and geometric coverage. Section IV-C describes how the Priority-driven Decoupled Sampler decouples semantic observation and geometric coverage and chooses a set of both geometric and semantic viewpoints. Subsequently, Section IV-D explains how the global planner optimizes the travel path through viewpoints. Finally, a Safe Aggressive Exploration State Machine is proposed in Section IV-E to enable efficient but safe exploration.
@@ IV-A Sensor Fusion and Pre-processing
Given the input measurement from both LiDAR and panoramic image, they are associated by:
@@
Using Eq. 1, 2, odometry, and object 2D detection, a raw point cloud is extracted for each instance mask of the object. Then, it is processed by applying DBSCAN clustering, selecting the closest group to the observation pose, and filtering out points lower than the ground. The instance mask, the associated instance point cloud, and the pose estimation at the current time step are packaged as a measurement $z$.
@@ IV-B Modular Semantic Target Mapping
Our plug-in semantic target mapping module builds a coarse-to-fine object-centric voxel model for each target object, updates the model constantly by real-time input measurement, and extracts a dense point cloud from the background SLAM algorithm.
@@ IV-B1 Data Association
Upon receiving a processed point cloud, the module matches it to existing object models using the bounding box intersection-over-union (IOU). If the best IOU falls below a threshold $T_{iou}$, a new object model is initialized.
@@ IV-B2 Model Update
The object model periodically incorporates new observations and processes multi-view observations statistically, accounting for potential false positives/negatives from varying view angles. Specifically, for each object model $O_{i}$, observations are grouped into $n$ observation sets. Each set corresponds to the observation space defined in Problem 2. Within each observation set, the model retains a keyframe measurement $z^{n}_{i}$ and stores its corresponding mask $m^{n}_{i}$ and pose estimation $p^{n}_{i}$. The key frame measurement $z^{n}_{i}$ is selected by the measurement with the highest mask score, defined as the following:
@@
where $V_{mask}$ and $V_{model}$ are bounding box volumes of the mask and current model, respectively. IOMax is the intersection over the maximal box. This score prioritizes masks giving more complete object measurement while avoiding taking an over-segmented point cloud with an unexpectedly large size.
However, false positive detection due to a complicated background could also produce a false positive keyframe measurement. To account for this, we model the validity of keyframes as a Bernoulli random variable. The optimal estimation of probability is given by the number of valid detections over the total number of observations in this observation set, where a valid observation means a mask score greater than a threshold $T_{score}$. Only keyframes with a true positive probability greater than $T_{p}$ will be kept. Each accepted observation is integrated into the voxel model, updating per-voxel probabilities. After this, for each observation set, the object model is ray-cast back to the key frame mask $m^{k}_{i}$ taking occlusions into account. The voxels outside the mask decrease its probability by $p_{dec}$, while those inside the mask increase the probability by $p_{inc}$. The above process is repeated for all observation sets.

### v2

# Semantic Exploration and Dense Mapping of Complex Environments using Ground Robot with Panoramic LiDAR-Camera Fusion

@@ Abstract
This paper presents a system for autonomous semantic exploration and dense semantic target mapping of a complex unknown environment using a ground robot equipped with a panoramic LiDAR-camera system. Existing approaches often struggle to strike a balance between collecting enough high-quality observations from multiple view angles and avoiding unnecessary repetitive traversal. To fill this gap, we propose a complete system that combines mapping and planning. We first redefine the task as completing both geometric coverage and semantic viewpoint observation. Subsequently, we manage semantic and geometric viewpoints separately and propose a novel Priority-driven Decoupled Local Sampler to generate local viewpoint sets. This allows for explicit multi-view semantic inspection and voxel coverage without unnecessary repetition. Building on this, we develop a hierarchical planner that ensures efficient global coverage. In addition, we propose a Safe Aggressive Exploration State Machine, which allows the robot to extend its exploration path planning into unknown areas while ensuring the robot’s safety with recovery behavior and adaptive sampling. Our system includes a modular semantic target mapping component designed to utilize odometry and point clouds from existing SLAM algorithms, enabling point-cloud-level dense semantic target mapping. We validate our approach through extensive experiments in both realistic simulations and complex real-world environments. Simulation results demonstrate that our planner achieves faster exploration and shorter travel distances while guaranteeing a specified number of multi-view inspections. Real-world experiments further confirm the effectiveness of the system in achieving accurate dense semantic object mapping of unstructured environments.
Index Terms:
Field Robots;Planning, Scheduling and Coordination; Engineering for Robotic Systems
@@ I Introduction
With the growing success of autonomous exploration algorithms, their applications are drawing increased attention from the research community. Semantic exploration is particularly important in scenarios such as factories and construction sites. Unlike Autonomous Aerial Vehicles (AAVs) commonly used in prior work, ground robots offer the advantage of accommodating more powerful sensors and computing systems, including high-resolution LiDAR and panoramic camera arrays. This paper focuses on semantic exploration and mapping for ground robots.
@@
Despite advancements, autonomous semantic exploration faces challenges. First, effective integration of mapping and planning is required, as the planner depends on semantic information to make informed decisions. Multi-view observation of target objects can help in many cases. For example, aggregating distinct observations around the object helps mitigate false detection results from the detector, and visual documentation like images from diverse perspectives also helps human supervisors in human-robot interaction scenarios. Therefore, the planner must actively seek informative viewpoints while computing a path for geometric coverage as short as possible. Existing methods either assign semantic weights to potential exploration goals that prioritize positions near target objects [1, 2] or sample viewpoints that cover unknown regions [3]. However, the former often causes repetitive movements around objects, while the latter neglects the importance of complete and diverse observation from multiple view angles. There is still a lack of a planner that strikes the balance for both.
Another difficulty is that robots operating in complicated environments need to balance exploration speed and collision risks. Existing planners usually restrict exploration to known free spaces to guarantee safety, but this over-conservative strategy results in short-sighted decision making and inefficient paths [4]. Although [5] extends global path guidance into unknown space, it still does not allow for the search and execution of local-level paths in unknown areas. However, executing aggressive exploration paths in unknown areas increases collision risks, especially at higher speeds, which limits planner performance.
@@
To address the above challenges, this paper proposes a system that performs semantic exploration planning and dense semantic target mapping in an unknown environment. We develop a modular, object-centric semantic target mapping component that operates on odometry and dense point cloud data from existing SLAM systems, without requiring tight integration with SLAM backend. The planning component defines the semantic exploration task as inspecting all unknown voxels and visiting evenly distributed semantic viewpoints to ensure high-quality mapping. To support the integration between mapping and planning, we introduce a hierarchical planner that manages semantic viewpoints and geometric exploration viewpoints in a unified framework. We propose a novel Priority-driven Decoupled Local Sampler to handle the two types of viewpoints separately and generate a minimum set of viewpoints that supports the completeness of both semantic and geometric coverage tasks. The sampler is then integrated with a global planner to generate the global optimal path by solving an asymmetric traveling salesman problem (ATSP). After that, to support aggressive exploration, a novel Safe Aggressive Exploration State Machine is proposed to ensure aggressive and resilient exploration in crowded regions.
@@
  • A modular semantic target mapping component based on panoramic LiDAR-camera fusion, directly leveraging the outputs from existing SLAM systems without requiring integration into back-end optimization.
@@
  • A safe aggressive exploration state machine that supports the robot in proactively selecting viewpoints and executing local paths in unexplored areas, while maintaining safety by actively checking and recovering.
  • Real-world tests in construction and industrial environments that reflect actual application deployment and validate the effectiveness of the complete system.
@@ II RELATED WORK
Autonomous robot exploration has made great progress over the decades [6, 7, 8, 9]. One stream of methods is based on frontiers, which are the boundaries between free space and unknown space. Yamauchi [6] greedily chooses the closest frontier as the next goal, while later methods [7, 10] apply more sophisticated frontier clustering, viewpoint generation, and path planning techniques. Another stream of methods is based on Next-Best-View (NBV) and graph sampling like RRT or PRM. In [11], a graph is built by sampling in safe free space and calculating the information gain for each node. The next best node with the highest average gain is chosen as the next goal. Later, Dang et al. [9] proposes rapidly-exploring random graphs for more rapid NBV exploration planning, and other methods [12, 13, 14, 15] continue to propose more carefully designed utility functions for information gain and graph structures. In recent years, more algorithms have been formulated as the traveling salesman problem (TSP) [7, 8, 16, 3] to improve the efficiency of traversing all viewpoints or graph nodes.
@@
At the same time, the SLAM community has an increasing interest in metric-semantic SLAM. Some of them use sparse representations such as cubes [17], ellipsoids [18], or hybrid geometric shapes [19], and combine the parameters into a factor graph to formulate an optimization problem. Others use dense representations such as meshes [20] or sufels [21].
@@
With the development of both areas, semantic exploration is gaining more attention [22, 1, 3, 23]. Tang et al. [2] define an object gain function to handle re-observation from multiple resolutions, and Papatheodorou et al. [1] propose multiple utility functions to include both background and object information gain to improve the quality of camera-based mesh reconstruction. Tao et al. [24] introduce semantic loop-closure viewpoints into exploration and formulate a Correlated Orienteering Problem (COP), but this formulation does not guarantee visiting all semantic viewpoints and requires a hard time-constraint estimation. Luo et al. [3] further propose a hierarchical planning framework to solve exploration with target search. VLM-based semantic navigation [23, 25] represents a recent sub-branch of exploration research, but recent work generally aim to find objects as soon as possible without pursuing full coverage of the environment. Meanwhile, the DARPA subterranean challenge facilitated semantic exploration research. Lei et al. [26] propose a pipeline to achieve high precision and recall of semantic detection and emphasize the importance of multi-view observation. Best et al. [27] utilized multi-sensor fusion during exploration and mapping.
@@
None of these methods, however, explicitly define multi-view observation viewpoints for object mapping in exploration, thus suffering from randomness due to unnecessary re-observation or inadequate observation. In addition, most algorithms are still based on the greedy NBV strategy with semantic weights, lacking careful consideration of global optimality. Our method fills this gap by explicitly defining the semantic viewpoint with a certain density around target objects while enabling efficient traversing considering both geometric coverage and semantic observation. Furthermore, we developed a Safe Aggressive State Machine to support aggressive exploration in complicated real-world environments.
@@ III PROBLEM FORMULATION
The goal of this paper is to enable a single ground robot, equipped with a panoramic LiDAR and camera fusion system, to autonomously plan its trajectory within an unknown environment and accurately map both the environment itself and specified target objects of interest. The robot should generate a dense point cloud for both the environment and the identified objects, while optimizing trajectory efficiency to ensure timely completion of the mapping tasks.
@@ III-A Robot and Sensor Model
[…]
@@ III-B Environment Model
The unknown environment is modeled as a bounded 3D volume $V\subset\mathbb{R}^{3}$, discretized into individual voxels $v$. Each voxel $v$ is associated with an occupancy probability $P_{o}(v)$. Before receiving any sensor measurements, each voxel is initialized as unknown with $P_{o}(v)=0.5$. Through the accumulation of sensor measurements and ray-casting, voxels are eventually classified as either occupied $V_{occ}$ or free $V_{free}$. Due to the constraints from the environment, robot and sensor configuration, some residual voxels $V_{res}$ may remain unmapped.
@@ III-C Problem Definition
Assume an unknown number $N\in\mathbf{N}$ of target objects, each associated with a semantic label $\lambda$ from a semantic label set $\mathbb{L}$. As discussed in Section I, the semantic exploration task must consider both target object observation and geometric coverage. We divide the space centering around a target object $O_{i}$ with radius $r$ into $M$ evenly distributed sectors. Each sector $S^{i}_{m}$ is defined as an observation space. Then, we define the semantic exploration task as follows:
Semantic Exploration: Given the sensor model and configuration $\psi_{robot}$, plan a path $\sigma_{robot}$ starting from an initial configuration $\xi_{robot}$ such that: (1) it covers at least one valid detection in every observation space $S^{i}_{m}$ for any target object $O_{i}$, whenever possible; (2) it observes all voxels so that $V_{free}\cup V_{occ}=V\setminus V_{res}$; and (3) it extracts all point cloud associated with object $O_{i}$ as well as the background.
@@ IV METHODOLOGY
The proposed system comprises a plug-in semantic mapping module and a hierarchical planner, as shown in Fig. 2. The mapping module, discussed in Section IV-B, takes as input the panoramic image $\mathbf{I}$, the LiDAR point cloud $\mathbf{P}$ and the estimation of existing odometry $\mathbf{O}$, subscribes the object bounding box $\mathbf{b_{o}}$ and the instance mask $\mathbf{m_{o}}$ from instance segmentation models such as [28], then outputs the dense point cloud of both the background environment $\mathbf{M_{b}}$ and target objects $\mathbf{M_{o}}$. Concurrently, the hierarchical planner generates the exploration path that ensures both semantic observation and geometric coverage. Section IV-C describes how the Priority-driven Decoupled Sampler decouples semantic observation and geometric coverage and selects geometric and semantic viewpoints. Subsequently, Section IV-D explains the global path calculation to travel through viewpoints. Finally, a Safe Aggressive Exploration State Machine is proposed in Section IV-E to enable efficient but safe exploration.
@@ IV-A Sensor Fusion and Pre-processing
The LiDAR-camera extrinsic calibration is derived from the known SPOT CAM and Ouster LiDAR CAD model. Time synchronization is handled by ROS’s approximate time policy with timestamps from the Boston Dynamics and Ouster API. Given the input measurement from both LiDAR and panoramic camera, they are associated by:
@@
Using Eq. 1, 2, odometry, object detection, and a raw point cloud are extracted for each instance mask. Then, it is processed by DBSCAN clustering, selecting the closest group to the observation pose, and filtering out points lower than the ground. The instance mask, the associated instance point cloud, and the pose estimation at the current time step are packaged as a measurement $z$.
@@ IV-B Modular Semantic Target Mapping
Our plug-in semantic target mapping module builds a coarse-to-fine object-centric voxel model for each target object, updates the model constantly with real-time input measurements, and extracts a dense point cloud from the background SLAM algorithm. It assumes a reasonably accurate odometry estimation. Although severe odometry drift could affect the mapping performance, the increasing availability of robust SLAM frameworks that provide the global pose graph optimization and loop-closure to reduce odometry drift could improve the system in future. Since it is out of scope of this paper, we will not discuss them here.
@@ IV-B1 Data Association
Upon receiving a processed point cloud, the module matches it to existing objects by bounding box intersection-over-union (IOU). If the best IOU falls below a threshold $T_{iou}$, a new object model is initialized.
@@ IV-B2 Model Update
The object model periodically incorporates new observations and processes multi-view observations statistically, accounting for potential false positives/negatives from varying view angles. Specifically, for each object model $O_{i}$, observations are grouped into $n$ observation sets. Each set corresponds to the observation space defined in Problem Definition. Within each observation set, the model retains a keyframe measurement $z^{n}_{i}$ and stores its corresponding mask $m^{n}_{i}$ and pose estimation $p^{n}_{i}$. The key frame is determined by choosing the best mask score:
@@
where $V_{mask}$ and $V_{model}$ are bounding box volumes of the mask and current model, respectively. IOMax is the intersection over the maximal box. This score prioritizes masks with complete object measurement while avoiding an over-segmented point cloud with an unexpectedly large size.
However, false positive detection due to a complicated background could also produce a false positive keyframe measurement. Therefore, we model the validity of keyframes as a Bernoulli random variable. The estimation of probability is given by the number of valid detections over the total number of observations in this observation set, where a valid observation means a mask score greater than a threshold $T_{score}$. Only keyframes with a true positive probability greater than $T_{p}$ will be kept. Each accepted observation is integrated into the voxel model, updating per-voxel probabilities. After this, for each observation set, the object model is ray-cast back to the key frame mask $m^{k}_{i}$, taking occlusions into account. The voxels outside the mask decrease its probability by $p_{dec}$, while those inside the mask increase the probability by $p_{inc}$. The above process is repeated for all observation sets.